# Top-of-atmosphere radiation fields over the last millennium reconstructed from proxies

### v1

@@ 4 Reconstruction over the last millennium
Our reconstruction consists of gridded, seasonal climate fields over 850–2000 CE. Results labeled as DA are the combined ensemble of the three single-model reconstructions. Annual means are taken from December to the following November. While we also reconstruct EEI directly, we found that the sum of its constituents -(RSR + OLR) has higher skill (Fig. S9a), and is thus shown as EEI here. Uncertainties refer to the very likely range (5th to 95th ensemble percentile, i.e., the 90% median credible interval).
@@ 4.1 Global means of temperature and radiation
The reconstructed global-mean SAT has a cooling trend over much of the last millennium, followed by warming over the 1900s (Fig. 5a). The SAT transition from a warm Medieval Climate Anomaly (MCA; c. 800–1200) into a colder Little Ice Age (LIA; c. 1300–1850) is a feature often found in last-millennium proxy reconstructions (e.g., Mann et al., 1999; Esper et al., 2002; Mann et al., 2009; PAGES 2k Consortium, 2019) and simulations (e.g., Fernández-Donado et al., 2013; Otto-Bliesner et al., 2016; Ljungqvist et al., 2019). Based on linear regression over 850–1850, we find a cooling trend of 0.30-0.30 K kyr-1 (see also Fig. 8a). There is also considerable multidecadal variability, some of which coincides with large volcanic eruptions; we will consider these events in greater detail in Section 44.5. The global-mean SST, OLR, and OHC300 are highly correlated with SAT (r >> 0.90; Fig. S8). The reconstruction is about 0.1 K warmer over 850–1850 than the multi-method ensemble from PAGES 2k Consortium (2019) but agrees within uncertainty.
@@
Figure 5: Global-mean time series over the last millennium. DA refers to the combined ensemble of the multi-model reconstructions. Shading denotes the 5th–95th percentile range. Bold lines have a 20-yr low-pass filter, thin gray lines are annual values. (a) SAT anomalies relative to 1961–1990, comparing to the multi-method ensemble and median from PAGES 2k Consortium (2019) and the HadCRUT5 instrumental dataset (Morice et al., 2021). Carets indicate volcanic eruptions, scaled by their volcanic stratospheric sulfur injection, from the eVolv2k (Toohey and Sigl, 2017) and the CMIP7 historical volcanic forcing datasets (Aubry and CMIP Climate Forcing Task Team, 2025; small eruptions with a VSSI below 0.5 Tg S removed). Those marked in orange are composited in Fig. 10. (b) Absolute energy imbalance. The DA EEI anomalies are made absolute by shifting to a zero-mean during 1600–1900, chosen to coincide with a period of relatively constant temperature. Colored lines show the rate of OHC change from the Gebbie and Huybers (2019) and Wu et al. (2025) reconstructions, expressed as equivalent EEI. The early 1100s and late 1200s volcanic clusters are annotated. (c) RSR and OLR. Atmosphere-only simulations (amip-hist; Zhou et al., 2016b), averaged over seven models, are shown in dashed lines as anomalies relative to 1961–1990. The DA anomalies in (c) are shifted to match the simulation 1871–1900 mean.
The absolute EEI governs changes in the energy budget, while we only reconstruct anomalies. We make the EEI anomalies absolute by shifting them to an 1600–1900 average of zero, chosen to coincide with relatively constant temperature and OHC, and thus no significant energy gain or loss over this period. Before 1600, the absolute EEI oscillates just below zero but has a slight positive trend (Figs. 5c and S9). This trend is due to a decrease in OLR, partially offset by a small increase in RSR. Note that this increase in EEI does not contradict the cooling temperature trend since the EEI remains negative. There is substantial interannual variability, occasionally on the order of 0.5 W m-2. Over the historical period, our results correlate well with the rate of change in 2000-m OHC from Wu et al. (2025). However, the multidecadal variability and the last-millennium trend show little in common with the rate of change in full-depth OHC from Gebbie and Huybers (2019, version OPT-0015), often having antiphased changes and differences as large as 1.0 W m-2. Further, the integrated energy imbalance, or energy inventory, disagrees with their OHC (Fig. S9). Note that the energy inventory is sensitive to the offset used to make the reconstructed EEI anomalies absolute, since any offset adds a linear function upon integration. Gebbie and Huybers (2019) obtain the OHC by subducting SSTs into the deep ocean. Their subducted SSTs, based on the Ocean2k proxy network (McGregor et al., 2015), are around 0.3 K colder in the pre-industrial period and show modern warming earlier than our SST anomalies (in the 1700s rather than the 1900s).
@@
The global-mean RSR and OLR are dominated by multidecadal variability, often associated with volcanic eruptions (Fig. 5c). The RSR is anticorrelated with SAT (annual r = 0.72-0.72), while the OLR has a strong positive correlation (annual r = 0.94). These correlations are about 50% stronger than in the last-millennium simulations, likely since we rely on temperature to reconstruct the radiation. In contrast, the correlation of EEI with SAT is weakly positive (annual r = 0.15), and similarly in the models. The ensemble spread for RSR is twice as large, mostly due to spread between the different model priors. The reconstructed RSR and OLR mostly agree with atmosphere-only simulations over the late 1800s, but then have opposite trends over the 1900s, similar to the PPE.
@@ 4.2 Context for recent energy imbalance
We consider the recent energy imbalance and its trend in the context of pre-industrial, natural variability. The EEI over 20-yr periods over 850–1850 in our reconstruction ranges from 0.36-0.36 to +0.27+0.27 (Fig. 6a). The distribution of this pre-industrial EEI provides context for the modern EEI, taken from CERES EBAF (Loeb et al., 2018), the 2025 update to the IPCC AR6 report (Forster et al., 2025; Smith et al., 2025), and Wu et al. (2025). In all datasets, the EEI over all 20-yr periods starting after 1970 exceeds any pre-industrial 20-yr-average EEI. Considering the 20th-century EEI from our reconstruction, this holds for all 20-yr periods starting after 1980.
The EEI trend over 24-yr periods over 850–1850 in our reconstruction ranges from 0.30-0.30 to +0.37+0.37 W m-2 dec-1 (Fig. 6b). The trend over the 24-yr CERES period (2001–2024) as observed from satellites and constrained by OHC observations in CERES EBAF is +0.45+0.45 W m-2 dec-1. Even a trend of +0.25+0.25 W m-2 dec-1, corresponding to the lower uncertainty bound of the CERES EBAF trend, is in the 97th percentile of pre-industrial variability.
Figure 6: Recent energy imbalance in the context of pre-industrial variability. (a) Absolute energy imbalance over 20-yr periods. The EEI reconstructed in this study is compared to CERES EBAF 4.2.1 data (Loeb et al., 2018) and to the 20-yr sliding window trend in the energy inventory reconstructions of Forster et al. (2025) and Wu et al. (2025). Shading denotes the 5th–95th percentile range. The histogram shows the distribution of 20-yr-mean, ensemble-mean EEI from our reconstruction over 850–1850, and dotted lines its minimum and maximum. The DA EEI anomalies used for the distribution are set to zero-mean during 1600–1900, chosen to coincide with a period of relatively constant temperature. (b) Energy imbalance trends over 24-yr periods, which is the number of full years in the CERES record. The histogram shows the distribution of 24-yr trends in the ensemble-mean EEI from our reconstruction over 850–1850. The trend from CERES EBAF is shown in blue, with an uncertainty of $\pm 0.20$ W m-2 dec-1 as in Raghuraman et al. (2021).
@@ 4.3 Sea ice
Sea ice influences high-latitude TOA radiation through surface albedo, lapse rate, and cloud feedbacks (Jenkins and Dai, 2021). Therefore, we reconstruct the SIC, from which the Arctic and Antarctic sea ice area (SIA) can be derived. While Arctic sea ice reconstructions over the last millennium exist (e.g., Brennan and Hakim, 2022; Meng et al., 2025), Antarctic sea ice has so far been elusive (Thomas et al., 2019).
SIA in both hemispheres increased with the cooling trend over the last millennium (Fig. 7). Arctic SIA agrees closely with Brennan and Hakim (2022), which is a proxy-based reconstruction using offline DA (no model forecasts). Over 850–1850, the annual-mean Arctic SIA grows by $0.5\times 10^{6}$ km2, mostly in the Barents Sea, while $0.9\times 10^{6}$ km2 are lost over 1850–2000 (Figs. S4 and S10). Antarctic SIA has less variability and more uncertainty when compared to Arctic SIA. Interannual to decadal variability in Antarctic SIA over 1700–2000 correlates highly with Dalaiden et al. (2023), another proxy-based reconstruction using offline DA. The annual-mean Antarctic SIA shrinks by $0.8\times 10^{6}$ km2 over 1850–2000. This shrinking is not present in Dalaiden et al. (2023)’s sea ice area, but similar to Cooper et al. (2025)’s reconstruction over 1900–2000 (Fig. S4).
During the growth period, the amplitude of the seasonal cycle in Arctic sea ice decreases, with a smaller boreal winter SIA and a larger summer SIA, peaking in JJA, relative to the annual mean. In contrast, the amplitude of the seasonal cycle in Antarctic sea ice increases during the last millennium, with peak growth in SON. The seasonal sea ice trends are synchronous with high-latitude insolation trends, which differ between the hemispheres (Fig. 7c,d). Extremes of insolation trends are smoothed in the seasonal averages, which obscures lead–lag relationships at monthly timescales.
Figure 7: Sea ice area over the last millennium for the Arctic and Antarctic. Anomalies are relative to 1961–1990, and shading denotes the 5th–95th percentile range. (a,b) Annual-mean SIA with 20-yr low-pass filter (bold lines). Arctic sea ice area is compared to Brennan and Hakim (2022)’s reconstruction. (c,d) Departures in seasonal trends from the annual-mean trend over 850–1850. The high-latitude (>> 60°) insolation trend (yellow) from Berger and Loutre (1991) is shown with an inverted y-axis.
@@ 4.4 Last-millennium trends
The global-mean cooling trend from the MCA into the LIA is a prominent feature in our reconstruction. Associated with it are regional variations, which we summarize here in terms of zonal-mean trends. Latitude–longitude gridded trends are shown in Fig. S11.
While all regions cool from the MCA into the LIA, polar amplification is evident in the SAT trend (Fig. 8a). Arctic amplification, measured by the ratio of the 67°N–90°N trend to the global-mean trend in SAT, is 3.9, similar to recent decades (Rantanen et al., 2022). Cooling is stronger in the Northern Hemisphere (NH; 0.37-0.37 K kyr-1) than in the Southern Hemisphere (SH; 0.22-0.22 K kyr-1). In comparison, global-mean temperature trends in the four CMIP6 last-millennium simulations range from 0.1-0.1 to +0.3+0.3 K kyr-1.
@@
Figure 8: Zonal-mean linear trends over 850–1850. Each panel shows (top) the annual mean and (bottom) the seasonal means. Yellow lines correspond to the CMIP6 last-millennium simulations from the MPI, CESM, MRI, and MIROC models. Numbers denote the global-mean trend (reconstruction in black with 5th–95th percentile range, models in yellow with min–max range). Shading denotes the 5th–95th percentile range.
@@
Trends in SST and OHC are closely related to each other (Fig. 8b,c), with cooling at all latitudes, particularly in the North Pacific, North Atlantic, and the tropical Pacific. However, there is significant uncertainty in the magnitude of reconstructed cooling in the northern mid-latitudes across model priors. Warming occurs around the Kuroshio–Oyashio Extension and the Gulf Stream (not shown).
@@
Trends in TOA radiation have a more complex spatial structure (Figs. 8d–f). The OLR trend is strongly negative and mostly confined to the tropics and subtropics, whereas the RSR trend is positive and extends globally. Together, this makes the global-mean EEI trend positive (toward energy gain). We find that the RSR and EEI trends are not significantly different from zero, and vary strongly by location and season (Figs. 9 and S11). The EEI trend is positive in the equatorial region and negative in mid-latitudes. The trend is most positive in MAM, driven by a tendency to gain energy in the equatorial east Pacific in both RSR and OLR. The opposite is the case in JJA, although the energy loss is offset by energy gain over the Maritime Continent.
Figure 9: Linear trend in EEI over 850–1850. (a) Annual-mean trend and (b–e) departures in seasonal trends from the annual-mean trend. The constituent RSR and OLR trends are shown in Fig. S11.

### v2

@@ 4 Reconstruction over the last millennium
Our reconstruction consists of gridded, seasonal climate fields over 850–2000 CE. Results labeled as DA are the combined ensemble of the five single-model reconstructions. Annual means are taken from December to the following November. Uncertainties refer to the very likely range (5th to 95th ensemble percentile, i.e., the 90% median credible interval).
@@ 4.1 Global means of temperature and radiation
The reconstructed global-mean SAT has a cooling trend over much of the last millennium, followed by warming over the 1900s (Fig. 5a). The SAT transition from a warm Medieval Climate Anomaly (MCA; c. 800–1200) into a colder Little Ice Age (LIA; c. 1300–1850) is a feature often found in last-millennium proxy reconstructions (e.g., Mann et al., 1999; Esper et al., 2002; Mann et al., 2009; PAGES 2k Consortium, 2019) and simulations (e.g., Fernández-Donado et al., 2013; Otto-Bliesner et al., 2016; Ljungqvist et al., 2019). Based on linear regression over 850–1850, we find a cooling trend of $0.18\,\mathrm{K\,kyr^{-1}}$ (see also Fig. 8a). There is also considerable multidecadal variability, some of which coincides with large volcanic eruptions; we will consider these events in greater detail in Section 44.5. The global-mean SST and OHC300 are highly correlated with SAT ($r>0.8$; OHC shown in Fig. 5d, SST in Fig. S11d). Our reconstruction has a similar mean value, multidecadal variability ($r=0.85$, $p\leq 0.01$ at 20-yr timescales), and trend as the multi-method ensemble from PAGES 2k Consortium (2019). This is also the case when comparing to LMR v2.1 (Tardif et al., 2019) and PHYDA (Steiger et al., 2018) in Fig. S11a. The NH extratropical land summer temperatures are at the cold end of comparable reconstructions (e.g., Wilson et al., 2016; Büntgen et al., 2021), and colder than instrumental temperatures before 1950, while the SH temperature agrees well with Neukom et al. (2014), as shown in Fig. S11b,c.
@@
Figure 5: Global-mean time series over the last millennium. DA refers to the combined ensemble of the multi-model reconstructions. Shading denotes the 5th–95th percentile range. Bold lines have a 20-yr low-pass filter; thin, faint lines are annual values. (a) SAT anomalies relative to 1961–1990, comparing to the multi-method ensemble from PAGES 2k Consortium (2019) and the HadCRUT5 instrumental dataset (Morice et al., 2021). (b) Absolute energy imbalance (rather than anomalies), representing the radiative response and natural forcing. The orange line shows the rate of OHC change from Gebbie and Huybers (2019), expressed as equivalent EEI and divided by 90%, which is the fraction of energy imbalance absorbed by the ocean (von Schuckmann et al., 2023). Gebbie and Huybers (2019) data after 1850 is dashed since it includes anthropogenic forcing, which we cannot reconstruct. Carets indicate volcanic eruptions, scaled by their volcanic stratospheric sulfur injection, from the eVolv2k (Toohey and Sigl, 2017) and the CMIP7 historical volcanic forcing datasets (Aubry et al., 2025; small eruptions with a VSSI below $0.5\,\mathrm{Tg\ S}$ removed). Those marked in orange are composited in Fig. 10. (c) RSR and OLR, constituting the EEI in (b). (d) Upper-ocean heat content anomalies, compared to the reconstructions by Gebbie and Huybers (2019) and Wu et al. (2025).
Figure 6: Recent energy imbalance in the context of pre-industrial variability. The EEI reconstructed in this study is compared to CERES EBAF Ed4.2.1 data (Doelling, 2025; Loeb et al., 2018), and to the 20-yr sliding window trend in the heat content reconstructions of IGCC (Forster et al., 2025; Smith et al., 2025) and Wu et al. (2025). The OHC from Wu et al. (2025) has been divided by 90%, which is the fraction of energy imbalance absorbed by the ocean (von Schuckmann et al., 2023). Shading denotes the 5th–95th percentile range, assuming independent errors when adding or subtracting time series. (a) Energy imbalance trends over 25-yr periods, which is the number of full years in the CERES record. The histogram shows the distribution of 25-yr trends in the ensemble-mean EEI from our reconstruction over 850–1850. The yellow line shows the kernel density estimate of 25-yr trends in the five CMIP6 last-millennium simulations. The trend from CERES EBAF is shown in blue, with an uncertainty of $\pm 0.20\,\mathrm{W\,m^{-2}\,dec^{-1}}$ as in Raghuraman et al. (2021). (b) Absolute energy imbalance over 20-yr periods. The black line is the sum of the reconstructed EEI and the anthropogenic forcing from IGCC. The vertical histogram shows the distribution of 20-yr-mean, ensemble-mean EEI from our reconstruction over 850–1850, and dotted lines show its minimum and maximum. (c) Response and natural forcing over 20-yr periods. The yellow line is the EEI from Wu et al. (2025) minus the anthropogenic forcing from IGCC, to which we anchor our EEI anomalies in order to obtain absolute values.
The absolute EEI governs changes in the energy budget (e.g., a positive EEI implies energy gain), whereas the reconstruction provides anomalies. To obtain absolute values, we anchor the reconstructed EEI anomalies, which contain the response and natural forcing, to an independent climatology of the total EEI minus the anthropogenic forcing, based on 20-yr running means with centers during 1880–1960. The equivalence follows from rearranging Eq. (2):
$\underbrace{R+F_{\text{nat}}}_{\text{DA}}=\underbrace{\text{EEI}}_{\text{Wu et al. (2025)}}-\underbrace{F_{\text{anthro}}}_{\text{IGCC}},$
where the EEI from Wu et al. (2025) is the rate of change of reconstructed ocean heat content, and the anthropogenic forcing from IGCC (Indicators of Global Climate Change, an update to metrics from the IPCC AR6 report; Forster et al., 2025; Smith et al., 2025) is estimated using empirical relationships. Our anomalies and the anchor timeseries have similar variability over the anchoring period of 1880–1960 (Fig. 6c).
The cooling trend over the last millennium coincides with energy loss (Fig. 5b), manifesting as a negative energy imbalance over 850–1500. After that, the energy imbalance is close to zero. The only other proxy-based estimate of the last-millennium energy budget is the OHC reconstruction from Gebbie and Huybers (2019, version OPT-0015). At multidecadal timescales, our reconstruction has a weakly negative correlation (r=0.10r=-0.10, $p=0.73$ at 50-yr timescales over 850–1850) with the Gebbie and Huybers (2019) estimate (Fig. S12a). The disagreement increases before 1100, although our reconstruction is likely less reliable during this early period, as discussed in Section 33.1. Over 1100–1850, our mean imbalance ($0.13\pm 0.13\,\mathrm{W\,m^{-2}}$) agrees well with theirs ($0.12\,\mathrm{W\,m^{-2}}$), and consequently, the integrated energy imbalance also agrees well (Fig. S12b), indicating a loss of about $1700\,\mathrm{ZJ}$ over 850–1850. After 1850, our reconstructions diverge because Gebbie and Huybers (2019)’s estimate also includes anthropogenic forcing. Further, their prescribed SSTs show stronger and earlier warming than ours (Figs. 5d and S11d).
@@
The global-mean RSR and OLR are dominated by multidecadal variability, often associated with volcanic eruptions (Fig. 5c). RSR is anticorrelated with SAT (annual r=0.78r=-0.78, $p\leq 0.01$), while OLR has a strong positive correlation (annual $r=0.97$, $p\leq 0.01$). These correlations between temperature and TOA radiation are about $20\%$ stronger than in last-millennium simulations, likely since our method depends on temperature to reconstruct radiation. In contrast, the correlation of EEI with SAT is weakly negative (annual r=0.21r=-0.21, $p=0.02$), while it is weakly positive in models. RSR has more variance than OLR at all timescales.
@@ 4.2 Context for recent energy imbalance
We consider the recent energy imbalance and its trend in the context of pre-industrial, natural variability. The EEI trend over 25-yr periods during 850–1850 in our reconstruction ranges from 0.26-0.26 to $0.32\,\mathrm{W\,m^{-2}\,dec^{-1}}$ (Fig. 6a). The distribution of these pre-industrial EEI trends provides context for the modern EEI trend. The trend over the 25-yr CERES satellite period (2001–2025) as provided in CERES EBAF data (Doelling, 2025; Loeb et al., 2018) is $0.44\,\mathrm{W\,m^{-2}\,dec^{-1}}$, far exceeding pre-industrial values. Even a trend of $0.24\,\mathrm{W\,m^{-2}\,dec^{-1}}$, corresponding to the lower uncertainty bound of the CERES EBAF trend, is in the 99th percentile of pre-industrial variability in our reconstruction. Last-millennium simulations have a slightly wider range of EEI variability, likely due to a stronger response to volcanic eruption. The CERES trend of $0.44\,\mathrm{W\,m^{-2}\,dec^{-1}}$ is in the 97th percentile of pre-industrial variability from last-millennium simulations.
EEI ranges from 0.41-0.41 to $0.07\,\mathrm{W\,m^{-2}}$ over 20-yr periods during 850–1850 in our ensemble-mean reconstruction (Fig. 6b). We compare this pre-industrial distribution to the modern EEI estimated from CERES EBAF satellite data, the IGCC heat inventory (Forster et al., 2025; Smith et al., 2025), and the OHC reconstruction by Wu et al. (2025). In these three modern datasets, the EEI over all 20-yr periods starting after 1915 exceeds any pre-industrial 20-yr-average EEI, although the EEI around 1960 almost falls within the pre-industrial distribution. Considering not just the distribution of the ensemble mean but also all ensemble members, the maximum pre-industrial EEI is $0.82\,\mathrm{W\,m^{-2}}$. With this more conservative threshold, the EEI for all 20-yr periods starting after 1990 is unprecedented in the pre-industrial period.
While our reconstruction does not contain the anthropogenic forcing, we can add its estimate from IGCC to our reconstructed EEI. The sum then agrees well with the total EEI from Wu et al. (2025) over 1880–1960, although there appears to be a 10-yr lag (Fig. 6b). After 1960, the response in our reconstruction is likely underestimated (Fig. 6c).
@@ 4.3 Sea ice
Sea ice influences high-latitude TOA radiation through surface albedo, lapse rate, and cloud feedbacks (Jenkins and Dai, 2021). We therefore evaluate Arctic and Antarctic sea-ice area (SIA), calculated as the sum of the products of SIC and grid-cell area in each hemisphere (e.g., Notz, 2014). While Arctic sea-ice reconstructions over the last millennium exist (e.g., Brennan and Hakim, 2022; Meng et al., 2025), Antarctic sea ice has so far been elusive (Thomas et al., 2019), and we caution that our Antarctic reconstruction shows much lower skill than the Arctic based on pseudoproxy experiments (Fig. S3) and instrumental validation (Figs. S8 and S9).
Figure 7: Annual-mean sea-ice area over the last millennium. Shading denotes the 5th–95th percentile range. We show anomalies relative to 1961–1990 with 20-yr low-pass filter and compare to the reconstructions from Brennan and Hakim (2022) for the Arctic and Dalaiden et al. (2023) for the Antarctic.
SIA in both hemispheres increased with the cooling trend over the last millennium (Fig. 7). Arctic SIA agrees closely with Brennan and Hakim (2022), which is a proxy-based reconstruction using offline DA (no model forecasts). Over 850–1850, the annual-mean Arctic SIA grows by $0.3\times 10^{6}\,\mathrm{km^{2}}$, mostly in the Barents Sea, while $0.9\times 10^{6}\,\mathrm{km^{2}}$ are lost over 1850–2000 (Figs. S8 and S13). The peak SIA during the LIA is slightly lower than in Brennan and Hakim (2022). Antarctic SIA has less variability and more uncertainty across model priors when compared to Arctic SIA. Interannual to decadal variability in Antarctic SIA over 1700–2000 correlates significantly with Dalaiden et al. (2023), another proxy-based reconstruction using offline DA. There is no clear trend in Antarctic sea-ice loss over 1850–2000 (Fig. S8) in our reconstruction, Cooper et al. (2025), and Dalaiden et al. (2023).
@@ 4.4 Last-millennium trends
The global-mean cooling trend from the MCA into the LIA is a prominent feature in our reconstruction. Associated with it are regional variations, which we summarize here in terms of zonal-mean trends, based on a linear regression over 850–1850; spatial trends are shown in Fig. S14.
@@
Figure 8: Zonal-mean trends over 850–1850. Each panel shows (top) the annual mean and (bottom) the seasonal means. Yellow lines correspond to the CMIP6 last-millennium simulations from the MPI, CESM, MRI, EC-Earth, and MIROC models. Numbers denote the global-mean trend (reconstruction in black with 5th–95th percentile range, models in yellow with min–max range). Shading denotes the 5th–95th percentile range.
Figure 9: Seasonal high-latitude (>> 60°) temperature trends over the last millennium for the Arctic and Antarctic. Trends are expressed as departures in seasonal temperature trends from the annual-mean trend over 850–1850. The high-latitude (>> 60°) insolation trend (yellow) from Berger and Loutre (1991) is shown as daily (dotted) and seasonal-mean (solid) values. Shading denotes the 5th–95th percentile range.
The global-mean cooling trend over 850–1850 is $0.18\,\mathrm{K\,kyr^{-1}}$, with a 5th–95th percentile range of 0.29-0.29 to $0.06\,\mathrm{K\,kyr^{-1}}$ (Fig. 8a). In the multi-method reconstruction from PAGES 2k Consortium (2019), the mean cooling trend is $0.30\,\mathrm{K\,kyr^{-1}}$, with a 5th–95th percentile range of 0.51-0.51 to $0.03\,\mathrm{K\,kyr^{-1}}$. In comparison, this trend in the five CMIP6 last-millennium simulations ranges from 0.11-0.11 to $0.28\,\mathrm{K\,kyr^{-1}}$. Coupled models thus tend to simulate last-millennium weak cooling or warming, contradicting proxy evidence. Cooling is stronger in the Northern Hemisphere (NH; $0.22\,\mathrm{K\,kyr^{-1}}$) than in the Southern Hemisphere (SH; $0.14\,\mathrm{K\,kyr^{-1}}$) in our reconstruction. While almost all regions cool from the MCA into the LIA, polar amplification is evident in the SAT trend. Arctic amplification, measured by the ratio of the 67°N–90°N trend to the global-mean trend in SAT, is around $4.4$, about $15\%$ higher than in recent decades (Rantanen et al., 2022).
@@
Trends in SST and OHC are closely related to each other (Fig. 8b,c), with cooling at all latitudes, particularly in the Pacific and North Atlantic. However, there is significant uncertainty in the magnitude of reconstructed cooling in the northern mid-latitudes across model priors. Warming occurs around the Kuroshio–Oyashio Extension and the Gulf Stream (Fig. S14b).
@@
Trends in TOA radiation have a more complex spatial structure that varies by season (Figs. 8d–f). The OLR trend is strongly negative and spatially defined by the Indo–Pacific convective region, whereas the RSR trend is positive and has large local trends at all latitudes. Together, this makes the global-mean EEI trend positive (toward energy gain), until the absolute EEI is approximately zero on centennial timescales after 1500. The RSR and EEI trends are not significantly different from zero, and vary strongly by location and season (Fig. S14d–f).
At high latitudes (>> 60°), seasonal differences in temperature trends are most pronounced (Fig. 9). The Arctic cooling trend is most enhanced in DJF and most subdued in JJA, which increases the amplitude of the seasonal cycle in Arctic temperatures during the last millennium. In contrast, Antarctic cooling is most enhanced in SON and most subdued in MAM and JJA, reducing the seasonal cycle amplitude. The seasonal temperature trends are compared with high-latitude insolation trends, which differ between the hemispheres. Extremes of insolation trends are smoothed in the seasonal averages, which obscures lead–lag relationships at seasonal timescales. However, a lag of up to one season is evident for the Arctic.